# LLM Analysis of 150+ years of German Parliamentary Debates on Migration Reveals Shift from Post-War Solidarity to Anti-Solidarity in the Last Decade

### v1

@@ 4.2.3 Distribution across parties
Figure 12 shows the distribution of solidarity and anti-solidarity subtypes across parties for the period 1949–1957. While solidarity was the dominant stance overall, parties varied in how they framed it, driven by ideological and strategic choices. Some parties used solidarity to promote national integration, while others employed it to defend regional concerns. Broadly, we divide parties into three categories: mainstream governing parties, expellee advocacy parties, and smaller parties with ideological or regional agendas.
@@
The main parties – governing CDU/CSU and FDP, and opposition SPD – supported the integration of expellees and refugees through legal and institutional measures, primarily expressing group-based or compassionate solidarity (see Figure 12). The CDU/CSU played a leading role in proposing the LAG, which framed solidarity as a national obligation toward fellow Germans, and even included expellee representatives in the party ranks hughes1999shouldering (49). The SPD also supported the law but voiced concerns about former Nazis and communist agents among the newcomers connor2018refugees (22), resulting in occasional anti-solidarity statements. The FDP, while formally aligned with CDU/CSU and SPD on key refugee issues, framed solidarity in more individualistic, exchange-based terms, opposing large-scale redistribution and favoring market-based solutions consistent with its liberal economic orientation connor2018refugees (22) (see Example 10 in Table 10, where FDP speakers argued that expellees should build their own homes through self-help and cooperative effort, with state support limited to loans and low-interest credit).
@@
Expellee advocacy parties (GB/BHE and DP) [^8] Gesamtdeutscher Block/Bund der Heimatvertriebenen und Entrechteten (GB/BHE) (en.: All-German Bloc/League of Expellees and Deprived of Rights) and Deutsche Partei (DP) (en.: German Party) explicitly positioned themselves as political representatives of displaced Germans. Following multiple successes in state elections, GB/BHE entered the federal government coalition in 1953, thereby exerting significant pressure on other parties to address the demands and concerns of expellees connor2018refugees (22). They framed solidarity primarily in terms of historical injustice and economic entitlement, calling for wealth redistribution, fair pensions, and affordable housing. Their rhetoric relied heavily on group-based solidarity, portraying expellees as full members of the national community who deserved compensation and full integration (see Example 11). Eventually, both major parties of the Federal Republic CDU and SPD managed to absorb the demands and concerns of the expellees, which made GB/BHE and DP less successful connor2018refugees (22, 54).
@@
Regionalist parties such as the WAV and BP [^9] Wirtschaftliche Aufbau-Vereinigung (en.: Economic Reconstruction Union) and Bayernpartei (en.: Bavarian Party) emphasized group-based solidarity, but restricted it to ethnic Germans. The WAV, a short-lived nationalist party, appealed to expellees through anti-communist rhetoric and demands to reclaim eastern territories kranenpohl1998zwischen (59) (Example 12). The BP, responding to Bavaria’s disproportionate refugee burden, called for collective responsibility in crisis, fairer redistribution and showed preference for groups like the Sudeten Germans braun2020101330 (16) (see Example 13,where the BP argued that Sudeten and Southeast Germans should be concentrated in Bavaria because they were ?closest to the Bavarian cultural sphere? and thus easier to integrate). Both parties also expressed the highest levels of anti-solidarity, either by excluding non-German migrants (WAV) connor2018refugees (22) or opposing specific groups such as Silesians (BP).
Figure 12: (Anti-)solidarity frames distribution across major parties in the Bundestag from 1949 to 1957, by election period. Percentages are normalized within each party, such that all subtype bars (solidarity and anti-solidarity) together sum to 100% per party. We include only those parties with a sufficient number of stanced instances $(N>100)$. The full label distribution across all parties is provided in Figure 6 in the Appendix.
The KPD [^10] Kommunistische Partei Deutschlands (en.: Communist Party of Germany) framed displaced populations through a lens of socialist integration, emphasizing group-based solidarity to support their incorporation into the new social order schwartz2008vertriebene (91). This solidarity, however, was assimilationist rather than inclusive: the party avoided the term Vertriebene, in line with Soviet repression of expellee identity heidemeyer2007deutsche (42), and this was reflected in its minimal use of empathic solidarity. The KPD referred almost exclusively to ?refugees?, advocating land reform, compensation, and integration as part of a broader effort to portray them as victims of fascism and capitalism and to gain political support connor2018refugees (22) (see Example 14, where the KPD argued that refugee employment could not be solved under capitalism).
Some parties active in the early postwar years, such as the KPD, WAV, and BP, did not stay in parliament for long, but they played an important role in shaping political debates of the time. Rather than merely reflecting existing social tensions, as suggested by lipset1967cleavage (67), these parties actively contributed to defining what solidarity meant and who was considered deserving of support. As suggested by thijssen2022s (98), such framings were shaped by the political and social challenges of the period and reflected broader efforts to maintain or renegotiate social cohesion.
@@ 4.3 Recent German history 2009-2025
Between 2009-2015, migration into Germany increased from just under 800,000 people per year to over 2 mio. people per year gessler2023politicization (41) in the context of the war in Syria. The peak in migration in 2015 has since been associated with the label ?migration crisis? gessler2023politicization (41, 36). Since then, the right-wing AfD has made electoral gains and entered the German federal parliament in 2017. Another spike in migration occurred in 2022 due to the Russian invasion of Ukraine bamf2025 (33). Alongside these historical events in the last four parliamentary periods (2009-2025), substantial shifts in solidarity occurred.
@@ 4.3.1 Overall solidarity distribution
Fig. 13 shows the (anti-)solidarity trends according to Llama-3.3-70b annotations between 2009-2025. We observe a decrease in solidarity between 2014-2018, a somewhat stable period between 2018-2022, and another decrease in solidarity 2022-2025. The decrease in solidarity is associated with an almost corresponding increase in anti-solidarity In terms of (anti-)solidarity subtypes, we observe a peak in compassionate solidarity in 2014, followed by a sharp decrease. The decrease in solidarity since 2022 corresponds to a decrease across solidarity subtypes, especially in compassionate and exchange-based solidarity. In terms of anti-solidarity, we observe a strong increase in group-based anti-solidarity, accounting for over 50% of all migrant-related statements that either expressed solidarity or anti-solidarity. Exchange-based anti-solidarity exhibited a small increase between 2015-2020 but dropped again with the beginning of the last parliamentary period in 2021.
Overall, the 2020s mark the first time in post-war history that the fraction of anti-solidarity statements exceeds that of solidarity statements. This key finding is not only visible in the data provided by Llama-3.3-70B but also in human annotations (Figure 2). To provide more insight into this finding, we contextualize it with recent Germany history and a deeper look into party distribution.
@@
Figure 13: Fig. 13(a) shows the fraction of solidarity, anti-solidarity, and mixed stance towards migrants from 2009 to 2025 (instances classified as ?None? are not shown, though included in the calculation). Fig. 13(b) shows the fraction of solidarity (left) and anti-solidarity (right) subtypes according to Llama-3.3-70B, where percentages are normalized within each year across all eight subtypes (thus, the curves in both panels together sum to 100% per year).
@@
The spike in compassionate solidarity 2014 coincides with a time of heightened public solidarity with Syrian refugees: 15.4% of German survey respondents indicated solidarity as the most important personal value in 2014, up from 12.3% in 2013 (DeLeur2023, 23, p. 38). Syrian refugees were commonly framed as needing help, motivating calls for a ?German welcome culture? Gebauer2023 (39) (see Example 15 in Table 14 in the Appendix). Over the course of 2015, a counter-narrative emerged, framing the influx of refugees as overburdening German capacities and as a ?refugee crisis? Gebauer2023 (39) (see Example 16). During the same period, the party AfD—founded in 2013 with a focus on EU criticism—aligned itself with anti-migrant protests and moved notably right, followed by substantial electoral gains in both state and national elections and the AfD entering the German parliament in 2017 Frey2020 (35, 39). Regarding the broader European context, DeLeur2023 (23) diagnose a ?solidarity fatigue? across Europe after 2015, driven by an inability to sustain prolonged compassion on the individual level and dwindling resources on the institutional level. The counter-narratives mentioned by Gebauer2023 (39, 35, 23) match several notions of anti-solidarity according to thijssen2022s (98): refugees as an outgroup that stands in competition to the German population (group-based anti-solidarity), refugees as taking away resources and not contributing sufficiently (exchange-based anti-solidarity), and refugees as not deserving of compassion (compassionate anti-solidarity); see Example 17 where refugees are described as not contributing, straining resources, and threatening the welfare of the ?real? German population. We also note a general shift towards right-wing sentiments in the general German population Zick2023 (108), which may be associated with heightened anti-solidarity.
@@
More recently, the Russian invasion in Ukraine and the subsequent migration of Ukrainian refugees into Germany triggered another spike in public solidarity, albeit much less prominent Weber2024 (104). In our data, solidarity increases slightly in 2022, followed by a continued, rapid decline. Bojarskich2025 (12) identified several factors promoting anti-solidarity towards Ukraine, such as conspiracy mentality and populist voting, but also fears about economic and societal consequences, indicating exchange-based (insufficient contribution) or group-based (competition between groups) anti-solidarity (Example 18). Additionally, anti-Ukrainian framing in parliamentary debates may have been reinforced by the AfD’s long-standing pro-Russian alignment arzheimer2023russia (5), including opposition to sanctions and calls to end German military aid to Ukraine.
In contrast to the postwar period, (anti-)solidarity shifts in the German parliament between 2009 and 2025 appear less attached to specific legislative initiatives. Because of Germany’s membership in the European Union, migration policy has become more complex and is negotiated between European institutions and member states geddes2020migration (40). Additionally, the AfD integrated anti-migrant talking points across a wide range of legislative topics. As such, we do not look deeper into specific laws passed during this time frame but rather focus on the trends for specific parties.
@@
Figure 14: Raw yearly counts of parliamentary statements by selected political parties, 2009–2025.
@@ 4.3.2 Distribution across parties
Figure 15: Distribution of solidarity and anti-solidarity subtypes over time (2009–2025) for selected German parties. For each party, percentages are normalized within each year across all eight subtypes (thus, the curves in both panels together sum to 100% per year). Note that AfD was first represented in the Bundestag in 2017, and the FDP was not represented from 2013 to 2017. Therefore, data points are absent for these parties during those respective periods.
Fig. 14 shows the raw numbers of migration-related statements per party and year, Fig. 15 shows the fine-grained solidarity trends.
@@
utters almost no expression of solidarity and the anti-solidarity subtypes remain relatively stable over time, with group-based anti-solidarity being consistently dominant (ca. 60% in 2018-2019 and ca. 75% in 2021-2025), although a notable number of samples with exchange-based solidarity (ca. 25%) occurs in the parliamentary period 2017-2021. Although the AfD is not the largest faction in parliament, they account for the largest share of anti-solidarity samples overall (72.5%). The sharp rise of anti-solidarity after 2017 (Figure 13(a)) aligns with the AfD’s entry into parliament and their rapid increase in migration-related utterances (Figure 14). This can be explained by the fact that many AfD speeches, across topics, contain references to migrants illegally crossing the border, criminality of migrants, or migrants imposing additional burden on the German society leaving less resources to address the actual topic of legislative debate (see Example 19 in Table 14 in the Appendix where an AfD member argues that refugees displace elderly people from care homes).
The center-right CDU/CSU and the liberal FDP
changed the most in our data. In the period 2009-2015, we observe a clear majority of solidarity samples, with a spike of compassionate solidarity for the CDU/CSU in 2014 (the FDP was not in parliament at that time). The main difference between both parties is more emphasis on exchange-based solidarity in the FDP, e.g. the contributions of incoming skilled workers to the German economy, motivating a new Fachkräfteeinwanderungsgesetz (en.: skilled labor migration law) in 2023 (see Example 20, where the FDP highlights the law as enabling foreign workers to contribute to the German economy). However, for both parties, we observe a rise in group-based anti-solidarity, accounting for 80% of all samples in 2025. When inspecting the samples, the framing of migration is mostly in legalistic terms, such as regulating and controlling migration and reducing numbers, marking migrants as an outgroup, which is classified as group-based anti-solidarity by Llama-3.3-70B (see Example 21 where a CDU/CSU speaker frames migration as a threat to political stability and an excessive burden on local infrastructure, with calls for deportations, faster procedures, and rejections at the border; and Example 22 where an FDP speaker portrays asylum as being misused for state-subsidized residence).
@@
The drastic shift toward group-based anti-solidarity over the last three years is insufficiently explained by recent literature. Partially, it may be that politicians from CDU/CSU and FDP take over populist rhetoric from AfD politicians who entered parliament since then Esguerra2023Accommodation (31). However,Lewandowsky2022 (65) and kesting2018discourse (53) did not observe effects of the AfD entering parliament on polarization, the use of populist language, or discourse quality. We also note that the first shift toward group-based anti-solidarity already occurred before the AfD enters parliament in 2017. Alternatively, it may be that CDU/CSU and FDP respond to an overall right-wing shift in the German population Zick2023 (108) by appealing to voters with anti-migration rhetoric and thereby providing a more moderate (but still anti-solidarity) alternative to the AfD. Overall, we cannot establish a full explanation but encourage further research into this striking and rapid trend.
@@
also exhibits a spike in compassionate solidarity in 2014 and a slow decline in solidarity towards more group-based anti-solidarity (ca. 20% in 2024). Part of this is explained by mis-classifications of Llama-3.3-70B because SPD speakers often cite anti-solidarity stances of other parties (especially AfD) and these citations are mis-classified as anti-solidarity by the speaker (see Example 23 where an SPD speaker cites AfD proposals putting foreigners under general suspicion in order to criticize them). However, there are also samples where SPD speakers refer positively to their own policies having enabled deportations of declined asylum seekers or the difficulties associated with managing migration, similar to the frames of CDU/CSU and FDP speakers (see Example 24).
@@
only display very few instances classified as anti-solidarity and almost all of them are misclassifications where Green or Linke speakers cite anti-solidarity statements or parliamentary motions by other parties (see Example 25 ). These samples are nonetheless interesting because they express a polarizing influence of the AfD entering parliament, leading left-wing parties to positioning themselves in contrast Atzpodien2022 (6). Overall, Die Grünen initially emphasized compassionate solidarity, but later shifted to group-based appeals stressing integration and rights. In the 2024/25 campaign, their rhetoric focused strongly on exchange-based arguments about migrants’ economic and societal contributions Gruene2025 (17) (see Example 26). Similarly, Die Linke began with predominantly compassionate appeals, framing migration as a humanitarian duty, but later emphasized migration as integral to society and advanced a detailed pro-immigration agenda, thereby increasing group-based solidarity. In the 2024/25 campaign, they moved further toward empathic solidarity, criticizing stereotyping and positioning themselves against exclusionary discourses DieLinke2021 (25, 26) (see Example 27). A special case is the BSW, which split off from the Linke party in 2024, partially due to explicit anti-migration positions. However, our data set contains only very few samples from this party (38 statements), so we cannot provide a reliable quantitative analysis.

### v2

@@ 4.2.3 Distribution across parties
Figure˜14 shows party distributions of solidarity and anti-solidarity in 1949–1957. Solidarity dominated overall, but parties framed it differently, shaped by ideology and strategy.
@@
The main parties – governing CDU/CSU and FDP, as well as opposition SPD – supported expellee and refugee integration through legal and institutional measures, primarily expressing compassionate or group-based solidarity (see Figure˜14). The CDU/CSU led in proposing the LAG, framing solidarity as a national duty toward fellow ethnic Germans, and even included expellee representatives in the party ranks (Hughes, 1999). The SPD also supported the law but voiced concerns about former Nazis and communist agents among the newcomers (Connor, 2018), resulting in occasional anti-solidarity statements. The FDP, while formally aligned with CDU/CSU and SPD on refugee issues, framed solidarity in more individualistic, exchange-based terms, opposing large-scale redistribution and favoring market-based solutions consistent with its liberal economic orientation (Connor, 2018) (see Example 17 in Table˜19).
Figure 14: Distribution of (anti-)solidarity subtypes across Bundestag parties, 1949–1957 (DSL-adjusted estimates; see definition (e) in Section˜4.2.2). Bars show each subtype’s share within a party, summing to 100%. The analysis includes parties with at least 150 DSL-adjusted stanced statements. Full label distributions appear in Table˜10 in the Appendix.
@@
The expellee advocacy party GB/BHE [^10] Gesamtdeutscher Block/Bund der Heimatvertriebenen und Entrechteten (GB/BHE) (en.: All-German Bloc/League of Expellees and Deprived of Rights). The DP (Deutsche Partei (en.: German Party)) also addressed expellee demands, especially in Lower Saxony because of their electoral weight there, but it began as a Heimatpartei (en.: ?homeland party?) rather than a refugee party (Aschoff, 2014). This is visible in the party’s share of group-based anti-solidarity, reflected in its defense of local interests (Example 19). positioned itself as a political representative of displaced Germans. Following multiple successes in state elections, GB/BHE, founded by and for expellees, entered the federal coalition in 1953 and pressured other parties to address expellee demands (Connor, 2018). It expressed compassionate and group-based solidarity, portraying expellees as full members of the national community entitled to compensation and integration (see Example 17). Additionally, the GB/BHE shows the highest share of empathic solidarity (15%), consistent with its role as an expellee party defining expellees as a distinct community. Over time, CDU and SPD managed to absorb major expellee demands, which made GB/BHE less successful (Connor, 2018; Kittel, 2020).
@@
Regionalist parties such as WAV and BP [^11] Wirtschaftliche Aufbau-Vereinigung (en.: Economic Reconstruction Union) and Bayernpartei (en.: Bavarian Party) emphasized compassionate solidarity, but restricted it to ethnic Germans. The short-lived WAV appealed to expellees through anti-communist rhetoric and demands to reclaim eastern territories (Kranenpohl, 1998) (Example 18). The BP, responding to Bavaria’s disproportionate refugee burden, called for invoked collective responsibility, fairer redistribution and favored groups like the Sudeten Germans (Braun and Dwenger, 2020), which corresponds to its high share of group-based anti-solidarity (see Example 18). Both parties also expressed the highest levels of mixed stance, either by excluding non-German migrants (WAV) (Connor, 2018) or opposing specific groups such as Silesians (BP).
The KPD [^12] Kommunistische Partei Deutschlands (en.: Communist Party of Germany) also addressed displaced populations, framing them through socialist integration and emphasizing compassionate and group-based solidarity, though in an assimilationist rather than inclusive sense. Avoiding the label Vertriebene in line with Soviet policy, the party instead spoke of ?refugees?, portraying them as victims of fascism and capitalism (Heidemeyer, 2007; Connor, 2018).
Parties like KPD, WAV, and BP did not stay in parliament for long, but nonetheless shaped early postwar debatesand helped define solidarity and its boundaries of deservingness. As suggested by Thijssen and Verheyen (2022), such framings were conditioned by the period’s political and social challenges and reflected broader efforts to maintain or renegotiate societal cohesion.
@@ 4.3 Recent German history 2009-2025
Between 2009 and 2015, annual migration to Germany rose from just under 800,000 to more than 2 million people, peaking in 2015 in the context of the Syrian war (Gessler and Hunger, 2023). This period has since been associated with the label ?migration crisis? (Gessler and Hunger, 2023; Fröhlich, 2025). Since then, the right-wing AfD has made electoral gains and entered the German federal parliament in 2017. Another wave of migration followed the 2022 Russian invasion of Ukraine (Federal Office for Migration and Refugees, 2025). Alongside these events in the last four parliamentary periods (2009-2025), substantial shifts in solidarity occurred.
@@ 4.3.1 Overall solidarity distribution
Figure 15: (Anti-)solidarity trends for the Migrant category, 2009–2025. Figure˜15(a): Yearly DSL-adjusted shares of solidarity, anti-solidarity, and mixed stance (see definition (c) in Section˜3.2.2). The none category is included in the normalization but omitted from the visualization. Figure˜15(b): Yearly DSL-adjusted shares of solidarity (left) and anti-solidarity (right) subtypes (see definition (c) in Section˜3.2.2). Within each year, subtype shares are renormalized across the eight solidarity and anti-solidarity categories (excluding mixed and none), summing to 100%.
Figure˜15 shows (anti-)solidarity trends from 2009–2025 (DSL-adjusted). Solidarity declined from 2014 to 2019, increased modestly between 2020 and 2022, and then saw a slight decline after 2022. This decrease in solidarity is associated with an almost corresponding increase in anti-solidarity. In terms of (anti-)solidarity subtypes, compassionate solidarity peaks around 2014–2016, drops sharply, briefly rebounds in 2021, [^13] This rebound coincides with the COVID-19 context, rising asylum applications, humanitarian responses to the Taliban takeover in Afghanistan, and the formation of the SPD–Greens–FDP coalition government (Brandt et al., 2024). and declines again. The decrease in solidarity since 2022 corresponds to a decrease across solidarity subtypes, especially in compassionate and exchange-based solidarity. In terms of anti-solidarity, group-based anti-solidarity rose strongly after 2017, accounting for over one third of all stanced migrant-related statements. Compassionate anti-solidarity spikes around 2020 and then declines after 2022.
Overall, the period since 2019 marks the first time in post-war history that the fraction of anti-solidarity statements exceeds that of solidarity statements in the German parliament. This key finding is visible both in the DSL-adjusted estimates but also in human annotations (Figure˜2). We next contextualize it with recent Germany history and a deeper look into party distribution.
@@
The 2014 spike in compassionate solidarity coincided with heightened public support for Syrian refugees: 15.4% of German survey respondents named solidarity their most important value in 2014, up from 12.3% in 2013 (De Leur et al., 2023, p. 38). Refugees were commonly framed as needing help, motivating calls for a ?German welcome culture? (Gebauer, 2023) (see Example 20 in Table˜23 in the Appendix). By early 2016, a counter-narrative emerged portraying refugees as overburdening German capacities and as contributing to a ?refugee crisis? (Gebauer, 2023) (see Example 20). During the same period, the party AfD—founded in 2013 with a focus on EU criticism—aligned itself with anti-migrant protests, shifted further towards the right of the political spectrum, and entered parliament in 2017 after major electoral gains (Frey, 2020; Gebauer, 2023). Regarding the broader European context, De Leur et al. (2023) diagnose a post-2016 ?solidarity fatigue?, driven by an inability to sustain prolonged compassion on the individual level and dwindling resources on the institutional level. These counter-narratives match several notions of anti-solidarity according to Thijssen and Verheyen (2022): refugees as an outgroup competing with German population (group-based, see Example 22), as undeserving of compassion (compassionate, see Example 23), and as straining resources without contributing (exchange-based, see Example 21). We also note a general shift towards right-wing sentiments in the general German population (Zick et al., 2023), which may be associated with heightened anti-solidarity.
@@
More recently, the Russian invasion of Ukraine and the resulting arrival of Ukrainian refugees in Germany, marking the largest net migration since 1950 (Atanisev et al., 2024), triggered another, though less prominent, spike in public solidarity (Weber et al., 2024a). In our data, solidarity increases slightly in 2022, followed by a continued decline. Bojarskich et al. (2025) identified several factors promoting anti-solidarity towards Ukraine among the German public, such as conspiracy mentality (i.e.  receptivity to contrarian narratives and disinformation emphasizing the risks of supporting Ukraine) and support for populist parties, but also fears about economic and societal consequences, indicating group-based (competition between groups) or exchange-based (insufficient contribution) anti-solidarity (Example 21). Anti-Ukrainian framing may also have been reinforced by the AfD’s pro-Russian stance, including opposition to sanctions and military aid (Arzheimer, 2023). In contrast to the postwar period, (anti-)solidarity shifts between 2009 and 2025 appear less tied to specific legislation, likely because much of German migration policy is negotiated at the EU level (Geddes et al., 2020). Additionally, the AfD integrated anti-migrant talking points across a wide range of legislative topics. As such, we focus less on legislation and more on party-specific trends.
@@
Figure 16: Raw yearly counts of parliamentary statements relating to migrants by selected political parties, 2009–2025.
@@ 4.3.2 Distribution across parties
Figure 17: DSL-adjusted yearly shares of (anti-)solidarity subtypes across Bundestag parties, 2009–2025 (see definition (c) in Section˜3.2.2). Within each year and party, subtype shares are renormalized across the eight solidarity and anti-solidarity categories so that they sum to 100%. Note: the AfD entered the Bundestag only in 2017, while the FDP was absent from 2013–2017; accordingly, no observations appear for these periods.
Figure˜16 shows the raw numbers of migration-related statements per party and year, Figure˜17 shows the fine-grained solidarity trends.
@@
expresses almost no solidarity, with anti-solidarity stable over time and dominated by group-based frames, fluctuating around 50% throughout 2017-2025, although a notable number of instances with compassionate and exchange-based anti-solidarity (ca. 25%) occurs, with the two frames alternating across years. Though one of the smallest factions in parliament, the AfD accounts for most anti-solidarity instances overall (53.3% over the whole period). The sharp rise of anti-solidarity after 2017 (Figure˜15(a)) aligns with the AfD’s entry into parliament and their rapid increase in migration-related utterances (Figure˜16). Many AfD speeches, across topics, link migrants to illegal border crossings, criminality, or additional burdens on the German society, without necessarily addressing the actual topic of legislative debate (see Example 21 in Table˜23 in the Appendix ).
The center-right CDU/CSU and liberal FDP
changed the most in our data. In 2009-2015, most statements expressed solidarity, with both parties focusing mostly on compassionate solidarity. However, the main difference between the two parties lies in their framing: the CDU/CSU emphasizes group-based solidarity, whereas the FDP foregrounds exchange-based solidarity, for instance by stressing the contributions of incoming skilled workers to the German economy, motivating a new Fachkräfteeinwanderungsgesetz (en.: Skilled Labor Migration Law) in 2023 (see Example 21). From the late 2010s onward, both parties shift toward anti-solidarity, most visibly in the CDU/CSU, where group-based anti-solidarity and compassionate anti-solidarity frames rise after the 2015 refugee crisis, decline during the Ukrainian refugee inflow in 2022, and resurge by 2025, reaching 33% and 24% of all frames, respectively. For the FDP, a comparable shift appears by 2025, where exchange-based, group-based, and compassionate anti-solidarity account for 16%, 12%, and 12%, respectively. When inspecting the instances, the framing of migration is mostly in legalistic terms, such as regulating and controlling migration and reducing numbers, marking migrants as an outgroup, for which group-based anti-solidarity has the highest DSL estimate among the anti-solidarity subtypes (see Example 22) and portraying migration as misuse of welfare benefits, corresponding to exchange-based anti-solidarity (see Example 22).
@@
The shift toward compassionate and group-based anti-solidarity over the last three years is insufficiently explained by recent literature. One possibility is that CDU/CSU and FDP adopted populist rhetoric from AfD after 2017 (Esguerra et al., 2023). However, Lewandowsky et al. (2022) and Kesting et al. (2018) did not observe effects of the AfD entering parliament on polarization, the use of populist language, or discourse quality. Notably, the first shift occurred even before the AfD entered parliament. Alternatively, it may be that CDU/CSU and FDP respond to an overall right-wing shift in the German population (Zick et al., 2023), appealing to voters with anti-migration rhetoric and thereby providing a more moderate (but still anti-solidarity) alternative to the AfD. Overall, our findings encourage further research into this striking and rapid trend.
@@
shows predominantly compassionate solidarity, with a spike in 2020, and is closely followed by group-based solidarity, which gains prominence post-2019, and almost taking over in 2025. This is consistent with the party’s humanitarian commitments and stronger emphasis on integration (Orhan, 2025). Additionally, exchange-based solidarity is present, at times reaching around 20%. In line with this perspective, SPD has repeatedly framed Germany as an Einwanderungsland (en.: country of immigration) and supported immigration policies aimed at attracting qualified workers to meet labor market demands (Sozialdemokratische Partei Deutschlands, 2021). However, there are also instances where SPD speakers express anti-solidarity by referring positively to their own policies having enabled deportations of declined asylum seekers or the difficulties associated with managing migration, similar to the frames of CDU/CSU and FDP speakers (see Example 22).
@@
only display very few instances classified as anti-solidarity. Overall, since 2021, Die Grünen moved from compassionate solidarity to group-based appeals, stressing integration and rights. In the 2024/25 campaign, and in the 2024/25 campaign strongly emphasized exchange-based arguments about migrants’ economic and social contributions (Bündnis 90/Die Grünen, 2025) (see Example 23). Die Linke likewise emphasized compassionate solidarity but since 2021 started increasingly framing migration as integral to society, advancing a detailed pro-immigration agenda and strengthening group-based solidarity. In 2024/25, the party continued to oppose stereotyping and position themselves against exclusionary discourses (Die Linke, 2021, 2025) (see Example 23). A special case is the BSW, which split from Linke party in 2024, partially due to explicit anti-migration positions. However, our data set contains only very few instances from this party (38 statements), so we cannot provide a reliable quantitative analysis.